OpenStudy (anonymous):

4. A ball is thrown into the air with an initial upward velocity of 60 ft/s. Its height (h) in feet after t seconds is given by the function h = –16t² + 60t + 6. What will the height be at t = 3 seconds?

OpenStudy (anonymous):

please vote http://www.signazon.com/contest/grad2014/lkkajqe5/

OpenStudy (anonymous):

lol i voted can i get any help

OpenStudy (anonymous):

@katragaddasaichandra

OpenStudy (anonymous):

u didnt vote

OpenStudy (anonymous):

i chcked

OpenStudy (anonymous):

check again dude

OpenStudy (anonymous):

i chckeed but u didnt

OpenStudy (anonymous):

did u type the words

OpenStudy (anonymous):

idid

OpenStudy (anonymous):

When the ball hits the ground, its position (height) will be zero. So, 0 = -16t² + 60t + 6 Solve for t, and t ≈ 3.8475, or about 4 seconds. The answer for 3.

OpenStudy (anonymous):

its asking what height its gonna be A.35ft B.40ft C.42ft D.45ft

OpenStudy (anonymous):

This does not factor, so we can do either of two things. One is to use the quadratic formula: t= [-B plus/minus the sqrt(B^2-4AC)]/(2A) In our case, A is -16, B is 60, and C=6 Using the quadratic formula, we get: t=[-60 plus/minus the sqrt(60^2-4*(-16)*6)]/(2*-16) t=[-60 plus/minus the sqrt(3984)]/-32 t=(-60+63.119)/-32 ~~~~ or ~~~~ t=(-60-63.119)/-32 t= -.097 ~~~~ or ~~~~ t=3.847 Since we cant have a negative time value, -.097 cannot be a solution. Thus, the ball hits the ground at t=3.847 seconds (rounded) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~... B) To find the heigh of the ball at t=3 seconds, we sumply plug 3 into our h(t)= -16t^2+60t+6 h(3)= -16*(3^2)+60*3+6 h(3)= -16*9+180+6 h(3)= 42 At t=3 seconds, the ball is 42 feet above the ground.

OpenStudy (anonymous):

it is c

OpenStudy (anonymous):

please vote

OpenStudy (anonymous):

http://www.signazon.com/contest/grad2014/lkkajqe5/

OpenStudy (anonymous):